OpenStudy (anonymous):

Solve by using the quadratic formula to solve x2 + 6x = -3. (1 points)

OpenStudy (mathstudent55):

First, add 3 to both sides of the equation. Then use the quadratic formula with a = 1, b = 6, and c = 3. Post what you get and ask questions if you have any.

OpenStudy (anonymous):

I go t -6 +- squareroot 24 over 2

OpenStudy (anonymous):

I did something wrong

OpenStudy (mathstudent55):

Good so far. Now simplify.

OpenStudy (anonymous):

when I simplified I got -6 +- sroot 12

OpenStudy (mathstudent55):

You have this so far, right? \( x = \dfrac{-6 \pm \sqrt{24}}{2} \)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

\( \sqrt{24} = \sqrt{4 \times 6} \) You can pull out \(\sqrt{4} = 2\), giving you \(2\sqrt{6} \).

OpenStudy (anonymous):

that isn't an option I got that also as I continued

OpenStudy (mathstudent55):

We're simplifying, We're not done yet. We'll get to the answer soon. \(x = \dfrac{-6 \pm \sqrt{24}}{2}\) \(x = \dfrac{-6 \pm 2\sqrt{6}}{2}\) \(x = \dfrac{(2)(-3) \pm (2)\sqrt{6}}{2}\) \(x = \dfrac{(2)[-3 \pm \sqrt{6}]}{2}\) \(x = -3 \pm \sqrt{6}\)

OpenStudy (anonymous):

thank you!!!

OpenStudy (mathstudent55):

wlcm